OpenStudy (anonymous):

A line has a slope of 2 and passes through the point (4, 0). What are the coordinates of another point that is on the line? a. (2, 0) b. (0, 2) c. (4, 6) d.(6, 4)

OpenStudy (anonymous):

someone please help.

OpenStudy (anonymous):

Ok, so we need to find the equation of the line. We have the slope and a point given to us. That means we need to use point-slope form. \[y-y_{1}=m(x-x _{1})\] Insert the point we have, and the slope. \[y-0=2(x-4)\] y-0=2x-8 y=2x-8-0 y=2x-8 Now graph the equation on a line, and see if any of the points are on the line.

OpenStudy (anonymous):

Here is a good tool to use: http://www.wolframalpha.com/input/?i=y%3D2x-8